# Thinking with Images for Multimodal Reasoning: Foundations, Methods, and Future Frontiers

### v2

@@ 5.3 Categories of Generative Thoughts
  • Interleaved Multimodal Reasoning. This paradigm represents a dynamic interplay between textual and visual thought, where reasoning is constructed through a dynamic, synergistic, and conversational sequence of generated text and images (Duan et al., 2025). The model’s thought process becomes a multimodal dialogue with itself: it might generate text to form a hypothesis, create an image to test it, and then generate more text to reflect on the visual evidence (Chern et al., 2025). This synergistic approach combines the logical structure of language with the grounded intuition of vision, enabling a powerful reasoning process for complex tasks like robotic planning and creative generation (Zhao et al., 2025).
@@ 5.4.1 SFT-Based Approaches
A significant thrust in this area is the development of unified models where SFT is crucial for balancing understanding and generation. Models like Chameleon (Team, 2024) and Emu2 (Sun et al., 2024a) are trained on vast quantities of interleaved mixed-modal data using a unified autoregressive objective to predict the next multimodal element. SFT is then applied via specialized “alignment recipes” or instruction tuning to hone performance on specific tasks. A diverse range of SFT strategies has emerged to manage the understanding-generation trade-off (Xie et al., 2024a; He et al., 2024a; Chen et al., 2024a). Some models adopt a sequential SFT approach, first training on understanding tasks and subsequently on generation, as seen with BLIP-3o (Chen et al., 2025b), which uses a curated instruction-tuning dataset for the generation phase. Others pursue efficiency by fusing pre-trained frozen components; GILL (Koh et al., 2023), for example, applies SFT only to lightweight mapping layers that connect a frozen LLM with image encoders/decoders. Further techniques include adjusting data ratios (Chen et al., 2025c), creating special visual tokens (Tong et al., 2024; Zheng et al., 2023), and freezing text-specific modules to preserve language capabilities while training new visual components (Shi et al., 2024a). Finally, models like BAGEL (Deng et al., 2025b) demonstrate that SFT on trillions of tokens from highly diverse, interleaved web data can unlock emergent capabilities like free-form image manipulation. These varied strategies underscore a core design challenge: carefully orchestrating SFT to create synergy, rather than destructive interference, between a model’s dual capabilities.
@@
Another direction uses SFT to enable models to generate visual thoughts as part of a transparent reasoning process. This approach moves beyond text-only chains of thought by externalizing the model’s visual and spatial reasoning. For instance, GoT (Fang et al., 2025) uses SFT to train a model to first produce an explicit language plan before generating an image that respects prescribed relationships. MVoT (Li et al., 2025a) fine-tunes models to think visually by generating image visualizations of their reasoning traces. Similarly, CoT-VLA (Zhao et al., 2025) prompts models to predict future image frames as visual subgoals before generating robotic actions. Others apply reasoning directly to the image creation pipeline. A more advanced paradigm, seen in Thinking with Generated Images (Chern et al., 2025), involves SFT on curated datasets of reasoning chains to enable models to spontaneously generate visual thinking steps like subgoals and self-critiques. This method of using SFT to create explicit multimodal reasoning chains that function as a visual scratchpad marks a shift towards more interpretable and robust AI.
@@
A clear trend is the convergence of techniques for generation and reasoning, with SFT acting as the unifying mechanism. Reasoning frameworks are increasingly applied to refine generation. For example, T2I-R1 (Jiang et al., 2025a) integrates a bi-level reasoning process into the text-to-image pipeline, using SFT and RL to improve output quality. Conversely, advanced generative mechanisms like diffusion models (Croitoru et al., 2023) are being integrated into reasoning frameworks. BLIP3-o (Chen et al., 2025b), for instance, uses SFT to train a diffusion transformer that generates rich image features as a basis for high-quality image synthesis. This intermingling demonstrates that strong generation requires sophisticated reasoning, while complex reasoning is enhanced by the ability to generate intermediate representations. SFT facilitates this powerful integration, blurring the lines between generative models and reasoning models and paving the way for more holistic AI systems.
@@ 5.4.2 RL-Based Approaches
The apex of thinking with images is Imagination. This paradigm is one where a model does not merely perceive or interact with existing visual information but instead generates entirely new visual content as intrinsic steps within its reasoning process. This represents a move from visual analysis to visual simulation and planning. The clearest embodiment of this concept is found in Visual Planning (Xu et al., 2025), which proposes a framework where models reason and plan purely through sequences of generated images, completely independent of text. This approach is trained with a novel Visual Planning via Reinforcement Learning framework. It demonstrates that a model can successfully navigate complex environments by imagining the sequence of future visual states, setting a benchmark for true visual-native reasoning.
@@
Effectively training these complex, imaginative processes requires sophisticated RL methodologies that go beyond simple correctness rewards. A key challenge is evaluating the quality of a generated image, which serves as an “imagined thought”. To address the reward modeling problem, recent work (Guo et al., 2025a) introduces PARM and PARM++, novel reward models designed specifically for autoregressive generation. They adaptively assess the generation process step-by-step for clarity and potential, with PARM++ even incorporating a reflection mechanism for self-correction, providing a more nuanced learning signal.
@@ 5.5 Conclusion and Future Frontiers
The emergence of intrinsic visual imagination represents the culmination of the “Thinking with Images” paradigm, marking a transition towards true cognitive autonomy. Instead of merely processing visual data, models are learning to use vision as an internal medium for thought. The three distinct paradigms discussed each advance this capability in a complementary way. Unified models provide a versatile and scalable foundation, efficiently balancing understanding and generation. Explicit reasoning via visual scratchpads introduces crucial transparency, making the model’s thought process interpretable and verifiable, which is invaluable for complex spatial problems. Finally, interleaved multimodal dialogue achieves a powerful synergy, combining the logical structure of language with the grounded intuition of vision in a dynamic reasoning loop. Collectively, these methods illustrate a clear trajectory where AI does not just consume visual content but genuinely thinks with it. This progress is transformative, yet the journey into artificial imagination is only beginning, opening up several critical frontiers for future research.
@@
The path for intrinsic visual imagination involves not only refining existing methods but also exploring fundamental questions about the nature of visual thought. The most promising frontiers aim to build more capable, efficient, and collaborative imaginative agents.
@@
  • Human-AI Collaborative Imagination. As AI develops more sophisticated imaginative capabilities, a transformative frontier will be human-AI co-imagination. This involves designing interfaces and interaction protocols that allow humans to guide, refine, and co-create with an AI’s visual thought process (Chen et al., 2025d). Such collaboration could unlock unprecedented creativity in fields from scientific discovery to artistic expression. This also brings forth critical ethical considerations regarding the responsible deployment of AI that can fabricate realistic visual content, necessitating robust safeguards and transparency.
@@ 6 Evaluations & Frameworks for Thinking with Images
Figure 4: Demonstration of the “Thinking with Images” paradigm across seven key evaluation domains. The figure showcases applications ranging from (a) Mathematical Reasoning and (b) STEM to (c) Perception Reasoning, (d) Code Generation, (e) Chart and Table Reasoning, (f) Real-world Applications, and (g) Puzzles and Games. In each domain, the model generates an intermediate visual thought to guide its reasoning process.
@@
Benchmark Characteristics Scenarios Eval Metrics
Setup
Mathematical Reasoning
MV-MATH (Wang et al., 2025f) Math reasoning in multi-visual contexts P R M G Auto
MathVista (Lu et al., 2024) Large-scale, multi-task visual math reasoning P R M G Hybrid
MM-MATH (Sun et al., 2024c) Middle school math with reasoning process evaluation P R G Hybrid
MathVision (Wang et al., 2024c) Competition-level multimodal math problems P R G Auto
MathVerse (Zhang et al., 2024d) Multi-version diagrammatic reasoning for robustness P R G Hybrid
OlympiadBench (He et al., 2024b) Olympiad-level math and physics problems P R G Hybrid
DynaMath (Zou et al., 2025) Dynamic benchmark for robustness evaluation P R G Auto
STEM
EMMA (Hao et al., 2025) Educational material analysis from textbooks P R M G Auto
SWE-bench MM (Yang et al., 2025b) Multimodal software engineering tasks P R G G Auto
MMMU (Yue et al., 2024a) Massive multimodal university-level subjects P R M Auto
MMMU-Pro (Yue et al., 2024b) Professional-level expert reasoning P R G M Auto
SFE (Zhou et al., 2025a) Expert-level scientific reasoning P R G Auto
PhysUniBench (Wang et al., 2025j) Undergraduate-level multimodal physics problems P R G M G Auto
Puzzles and Games
INGVP (Zhang et al., 2024e) Spatial imagination & rule-following puzzles P R M Auto
VGRP-Bench (Ren et al., 2025) Grid-based puzzles with logical reasoning P R G Auto
PuzzleBench (Zhang et al., 2025c) Visual QA based on puzzles P R G Auto
EnigmaEVAL (Wang et al., 2025k) Multi-step reasoning on unstructured puzzles from competitions P R G G Auto
PuzzleWorld (Li et al., 2025h) Large-scale benchmark for creative open-ended puzzle solving P R G G Hybrid
VCBench (Li et al., 2024b) Visual commonsense, abstraction, and control P R G LLM
VSP (Wu et al., 2024f) Perception and reasoning in spatial planning P R G G Auto
Perception Reasoning
V*Bench (Wu and Xie, 2023) Fine-grained Perception on High-Resolution Images P R M Auto
BLINK (Huang et al., 2023) Fine-grained Visual Perception P M Hybrid
Verify (Kumar et al., 2024) Visual Explanation and Multi-step Logical Reasoning P R G LLM
CogBench (Zheng et al., 2024c) Assessing Cognitive Abilities of LMMs P R G LLM
VBench (Wu et al., 2023b) Comprehensive Video Generation Evaluation G G Hybrid
Code Generation
MMCod (Li et al., 2024c) Visual Content-Grounded Code Generation P R G G Auto
Plot2Code (Wu et al., 2025b) Mathematical Plot-to-Code Generation P G G Hybrid
Design2Code (Si et al., 2025) Design Layout-to-Webpage Generation P G G Auto
Sketch2Code (Li et al., 2025i) Hand-drawn Sketch-to-Code Generation P G G Hybrid
WebSight (Laurençon et al., 2024) Web Screenshot-to-Code generation P G G Hybrid
Chart and Table Reasoning
ChartX (Xia et al., 2024b) Chart Type Recognition and Data Extraction P R G M G Hybrid
CharXiv (Wang et al., 2024d) Scientific Chart Understanding and QA P R M G Hybrid
M3SciQA (Li et al., 2024d) Multimodal Scientific Chart and Table QA P R M Hybrid
ChartSumm (Tang et al., 2025) Chart Summarization P R G G LLM
MultiChartQA (Zhu et al., 2025c) Multi-chart Question Answering and Reasoning P R M G Hybrid
Ocr-Reasoning (Huang et al., 2025b) Reasoning on OCR-rich Images P R G Hybrid
Real-world Applications
MULTI (Zhu et al., 2024) Diverse Real-World Multimodal Tasks P R G Hybrid
VisNumBench (Weng et al., 2025b) Visual Numerical Reasoning and Counting P R M G Auto
MME-Realworld (Fu et al., 2023) Holistic Evaluation of Perception and Cognition P R G Auto
MEGA-BENCH (Chen et al., 2024b) Large-scale suite with diverse real-world tasks P R G G Hybrid
Table 4: Further benchmarks assessing the “Thinking with Images” paradigm. The columns detail each benchmark’s Characteristics; the core cognitive Scenarios tested (P: Active Perception, R: Visual Reasoning, G: Generative Planning); the Eval Setup (M: multi-choice, G: generation); and the Metrics (Auto: automatic metrics, LLM: LLM-as-a-Judge, Hybrid: a combination of both).
@@ 6.1 Evaluations Benchmarks
The benchmarks for “Thinking with Images” measure a model’s ability to perform complex tasks that extend beyond static visual perception. As illustrated by the diverse examples in Figure 4, these tasks span multiple reasoning domains, from mathematical problem-solving to real-world applications. To provide a structured overview of this critical landscape, we organize these benchmarks into seven primary categories based on their core objectives. These categories are summarized in Table 4, with detailed benchmarks presented in Table 5 and Table 6 in Appendix.
@@
Multimodal mathematical problems are designed to evaluate the capabilities of LMMs in complex visual perception, multi-step reasoning, and the integration of mathematical knowledge. In this domain, substantial attention has been devoted to geometry problems (Cao and Xiao, 2022; Chen et al., 2022a; Fu et al., 2025b; Zhang et al., 2024f), which require interpreting fundamental visual elements and performing multi-step logical reasoning. Subsequently, the evaluation scope has broadened to support more comprehensive assessments (Lu et al., 2024; Sun et al., 2024c; Wang et al., 2025f; Sun et al., 2025), resulting in benchmarks that span a wide range of mathematical tasks, including algebra, analytic geometry, and arithmetic. More recently, the emergence of highly capable reasoning LMMs (Guo et al., 2025b; Gemini Team, Google, 2025) has spurred the development of a new generation of benchmarks aimed at advancing the frontiers of AI reasoning. These datasets present competition-level and highly complex problems, often drawn from mathematical olympiads (Wang et al., 2024c; He et al., 2024b; Gupta et al., 2024a), to rigorously evaluate the limits of current systems. In parallel, specialized evaluations have emerged to probe specific dimensions such as model robustness to diagrammatic variations (Zhang et al., 2024d; Zou et al., 2025), reasoning under minimal or no textual context (Kamoi et al., 2024; Xiao et al., 2024), and the ability to generate structured outputs such as visual diagrams or natural language explanations (Li et al., 2025b; Park et al., 2025). The core challenge in these tasks is that the solution often requires constructing intermediate visual constructs that is not present in the initial input, such as adding an auxiliary line or performing a geometric transformation. This setting uniquely reveals a model’s capacity for constructive visual reasoning. It reflects the ability to treat an image not as a static scene to be described, but as a dynamic workspace for logical construction and problem solving, which is an essential component of the “Thinking with Images” paradigm.
@@
STEM (Science, Technology, Engineering, and Mathematics) benchmarks are designed to evaluate the problem-solving capabilities of LMMs across scientific and technical domains. Early efforts established broad benchmarks in science education, targeting cross-disciplinary reasoning at the K–12 level (Lu et al., 2022; Hao et al., 2025; Zhou et al., 2025b; Ye et al., 2025). Subsequent work expanded to university-level assessments requiring deep subject knowledge (Yue et al., 2024a; Xia et al., 2025b; Wang et al., 2025j), and more recently, to professional-level benchmarks designed to evaluate expert-level reasoning in advanced systems (Yue et al., 2024b; Zhou et al., 2025a). In parallel, specialized benchmarks have been developed to assess distinct skills such as chart interpretation (Jian et al., 2025), isomorphic problem solving (Fu et al., 2024a), and multimodal software engineering (Yang et al., 2025b). Unlike the constructive but often static reasoning observed in mathematical tasks, STEM problems demand dynamic reasoning over temporal and causal phenomena, including physical motion, chemical reactions, and mechanical operations. This shift enhances the “Thinking with Images” paradigm by redefining the image as a dynamic representation of real-world processes. As a result, these benchmarks provide a testbed for evaluating a model’s understanding of physical laws and, more broadly, the sophistication of its emerging internal world model.
@@
Benchmarks for puzzles and games are designed to evaluate the capacity of LMMs for strategic planning and reasoning in dynamic, rule-governed environments. Early benchmarks introduced a range of tasks, from visual puzzles assessing logical and spatial reasoning (Zhang et al., 2024e; Ren et al., 2025; Zhang et al., 2025c; Wu et al., 2024f) to interactive games that require understanding of game mechanics and state transitions (Wang et al., 2025l; Tong et al., 2025a; Paglieri et al., 2024). More recent benchmarks have increased in complexity, incorporating dynamic video-based contexts that test physical and commonsense reasoning (Zhang et al., 2025g; Cao et al., 2024; Li et al., 2024b), as well as unstructured, competition-style puzzles demanding creative and open-ended problem solving (Wang et al., 2025k; Li et al., 2025h). Distinct from reasoning in STEM domains, the core challenge of games lies in long-horizon, prospective reasoning. This represents a form of “Thinking with Images” in which a model must visually simulate complex action sequences on an interactive game board in order to devise an optimal plan. The scalability and structured nature of these environments also makes them suitable for both training and evaluating the planning and reasoning capabilities of LMMs.
@@
Perception and reasoning benchmarks evaluate the visual capabilities of LMMs, encompassing the accurate interpretation of fine-grained visual elements and multi-step visual reasoning. One line of these tasks focuses on perceptual acuity, using high-resolution and detail-rich imagery to assess fine-grained visual recognition (Wu and Xie, 2023; Huang et al., 2023; Wang et al., 2024e). Another line targets complex cognitive functions, including multi-step reasoning (Kumar et al., 2024), object hallucination detection (Li et al., 2023a, 2024e), and cognitive evaluations (Zheng et al., 2024c; Gupta et al., 2024b; Wu et al., 2023b). These benchmarks assess whether a model can move beyond superficial pattern recognition to develop a deeper, structured understanding of visual scenes that supports robust reasoning. This capability establishes the perceptual and cognitive foundation necessary for the “Thinking with Images” paradigm.
@@
Code generation from visual inputs constitutes an application domain for the “Thinking with Images” paradigm, requiring LMMs to translate complex visual specifications into executable programming logic. Benchmarks in this area evaluate this ability across diverse contexts, such as interpreting mathematical plots and general programming-related diagrams (Li et al., 2024c; Wu et al., 2025b). One significant application is front-end development, where models are tasked with converting visual designs, ranging from hand-drawn sketches to polished UI screenshots, into functional webpage code (Si et al., 2025; Li et al., 2025i; Laurençon et al., 2024). This visual-to-code process exemplifies a distinctive form of “Thinking with Images.” It requires the model to function as a visual decompiler, parsing spatial layouts and design hierarchies to reconstruct a visual artifact into its underlying structural logic. The ability to convert visual intent into precise, executable code represents a practically valuable application of advanced reasoning LMMs.
@@
Real-world application benchmarks evaluate the ability of LMMs to handle diverse, authentic tasks that reflect practical scenarios from daily-life contexts. Work involves comprehensive, multi-task benchmarks that use authentic materials like examination questions to assess broad competency across a wide range of subjects (Zhu et al., 2024; Zhang et al., 2023c; Fu et al., 2023; Chen et al., 2024b). In parallel, specialized benchmarks probe performance on specific daily-life skills, such as temporal reasoning with clocks and calendars, or visual numerical reasoning and counting (Weng et al., 2025b). The core challenge of these benchmarks is to apply reasoning to noisy, unstructured visual data by leveraging broad commonsense and world knowledge. This serves as the ultimate test for the “Thinking with Images” paradigm, assessing whether a model’s visual reasoning capabilities can generalize from controlled settings to practical, real-world utility.

### v3

@@ 5.3 Categories of Generative Thoughts
  • Interleaved Multimodal Reasoning. This paradigm represents a dynamic interplay between textual and visual thought, where reasoning is constructed through a dynamic, synergistic, and conversational sequence of generated text and images (Duan et al., 2025). The model’s thought process becomes a multimodal dialogue with itself: it might generate text to form a hypothesis, create an image to test it, and then generate more text to reflect on the visual evidence (Chern et al., 2025). This synergistic approach combines the logical structure of language with the grounded intuition of vision, enabling a powerful reasoning process for complex tasks like robotic planning and creative generation (Zhao et al., 2025a).
@@ 5.4.1 SFT-Based Approaches
A significant push in this area is the development of unified models where SFT is crucial for balancing understanding and generation. Models like Chameleon (Team, 2024) and Emu2 (Sun et al., 2024a) are trained on vast quantities of interleaved mixed-modal data using a unified autoregressive objective to predict the next multimodal element. SFT is then applied via specialized alignment recipes or instruction tuning to hone performance on specific tasks. A diverse range of SFT strategies has emerged to manage the understanding-generation trade-off (Xie et al., 2024a; He et al., 2024a; Chen et al., 2024a). Some models adopt a sequential SFT approach, first training on understanding tasks and subsequently on generation, as seen with BLIP-3o (Chen et al., 2025b), which uses a curated instruction-tuning dataset for the generation phase. Others pursue efficiency by fusing pre-trained frozen components; GILL (Koh et al., 2023), for example, applies SFT only to lightweight mapping layers that connect a frozen LLM with image encoders/decoders. Further techniques include adjusting data ratios (Chen et al., 2025c), creating special visual tokens (Tong et al., 2024; Zheng et al., 2023), and freezing text-specific modules to preserve language capabilities while training new visual components (Shi et al., 2024a). Finally, models like BAGEL (Deng et al., 2025b) demonstrate that SFT on trillions of tokens from highly diverse, interleaved web data can unlock emergent capabilities like free-form image manipulation. These varied strategies highlight a core design challenge: carefully orchestrating SFT to create synergy, rather than destructive interference, between a model’s dual capabilities.
@@
Another direction uses SFT to enable models to generate visual thoughts as part of a transparent reasoning process. This approach moves beyond text-only chains of thought by externalizing the model’s visual and spatial reasoning. For instance, GoT (Fang et al., 2025) uses SFT to train a model to first produce an explicit language plan before generating an image that respects prescribed relationships. MVoT (Li et al., 2025a) fine-tunes models to think visually by generating image visualizations of their reasoning traces. Similarly, CoT-VLA (Zhao et al., 2025a) prompts models to predict future image frames as visual subgoals before generating robotic actions. Others apply reasoning directly to the image creation pipeline. A more advanced paradigm, seen in Thinking with Generated Images (Chern et al., 2025), involves SFT on curated datasets of reasoning chains to enable models to spontaneously generate visual thinking steps like subgoals and self-critiques. This method of using SFT to create explicit multimodal reasoning chains that function as a visual scratchpad marks a shift towards more interpretable and robust AI.
@@
A clear trend is the convergence of techniques for generation and reasoning, with SFT acting as the unifying mechanism. Reasoning frameworks are increasingly applied to refine generation. For example, T2I-R1 (Jiang et al., 2025a) integrates a bi-level reasoning process into the text-to-image pipeline, using SFT and RL to improve output quality. Conversely, advanced generative mechanisms like diffusion models (Croitoru et al., 2023) are being integrated into reasoning frameworks. BLIP3-o (Chen et al., 2025b), for instance, uses SFT to train a diffusion transformer that generates rich image features as a basis for high-quality image synthesis. This interplay shows that robust generation demands advanced reasoning, while powerful reasoning is amplified by the ability to generate intermediate steps. SFT facilitates this powerful integration, blurring the lines between generative models and reasoning models and paving the way for more holistic AI systems.
@@ 5.4.2 RL-Based Approaches
One of the most advanced forms of thinking with images is imagination, in which a model not only perceive or interact with existing visual information but also generates entirely new visual content as intrinsic steps within its reasoning process, representing a move from visual analysis to visual simulation and planning. The clearest embodiment of this concept is found in Visual Planning (Xu et al., 2025), which proposes a framework where models reason and plan purely through sequences of generated images, completely independent of text. This approach is trained with a novel Visual Planning via Reinforcement Learning framework. It demonstrates that a model can successfully navigate complex environments by imagining the sequence of future visual states, setting a benchmark for true visual-native reasoning.
@@
Effectively training these complex, imaginative processes requires advanced RL methodologies that go beyond simple correctness rewards. A key challenge is evaluating the quality of a generated image, which serves as an “imagined thought”. To address the reward modeling problem, recent work (Guo et al., 2025a) introduces PARM and PARM++, novel reward models designed specifically for autoregressive generation. They adaptively assess the generation process step-by-step for clarity and potential, with PARM++ even incorporating a reflection mechanism for self-correction, providing a more in-depth learning signal.
@@ 5.5 Conclusion and Future Frontiers
The emergence of intrinsic visual imagination represents the culmination of the “Thinking with Images” paradigm, marking a transition towards true cognitive autonomy. Instead of merely processing visual data, models are learning to use vision as an internal medium for thought. The three distinct paradigms discussed each advance this capability in a complementary way. Unified models provide a versatile and scalable foundation, efficiently balancing understanding and generation. Explicit reasoning via visual scratchpads introduces crucial transparency, making the model’s thought process interpretable and verifiable, which is invaluable for complex spatial problems. Finally, interleaved multimodal dialogue achieves a powerful synergy, combining the logical structure of language with the grounded intuition of vision in a dynamic reasoning loop. Collectively, these methods chart a clear trajectory for AI, moving it beyond merely consuming visual content to genuinely thinking with it.
@@
While this progress is transformative, the journey into artificial imagination has only just begun, opening up several critical frontiers for future research. The path for intrinsic visual imagination involves not only refining existing methods but also exploring fundamental questions about the nature of visual thought. The most promising frontiers aim to build more capable, efficient, and collaborative imaginative agents.
@@
  • Human-AI Collaborative Imagination. As AI develops more advanced imaginative capabilities, a transformative frontier will be human-AI co-imagination. This involves designing interfaces and interaction protocols that allow humans to guide, refine, and co-create with an AI’s visual thought process (Chen et al., 2025d). Such collaboration could unlock unprecedented creativity in fields from scientific discovery to artistic expression. This also brings forth critical ethical considerations regarding the responsible deployment of AI that can fabricate realistic visual content, necessitating robust safeguards and transparency.
@@ 6 Evaluations & Frameworks for Thinking with Images
Figure 4: Demonstration of the “Thinking with Images” paradigm across seven key evaluation domains. The figure showcases applications ranging from (a) Mathematical Reasoning and (b) STEM to (c) Perception Reasoning, (d) Code Generation, (e) Chart and Table Reasoning, (f) Real-world Applications, and (g) Puzzles and Games.
@@ 6.1 Evaluations Benchmarks
The benchmarks for “Thinking with Images” measure a model’s ability to perform complex tasks that extend beyond static visual perception. As illustrated by the diverse examples in Figure 4, these tasks span multiple reasoning domains, from mathematical problem-solving to real-world applications. To provide a structured overview of this critical landscape, we organize these benchmarks into seven primary categories based on their core objectives.
@@
Multimodal mathematical problems are designed to evaluate the capabilities of LMMs in complex visual perception, multi-step reasoning, and the integration of mathematical knowledge. Within this domain, geometry problems represent a key challenge in this domain (Cao and Xiao, 2022; Chen et al., 2022a; Fu et al., 2025b; Zhang et al., 2024d), as they require LMMs to perceive fundamental visual elements and perform multi-step logical reasoning, thereby serving as an effective benchmark for evaluating both the visual and linguistic reasoning capabilities of LMMs. To further assess the reasoning capabilities of LMMs across a broader range of multimodal mathematical scenarios beyond geometry, recent studies have expanded the evaluation scope (Lu et al., 2024; Sun et al., 2024c; Wang et al., 2025f; Sun et al., 2025) to encompass diverse mathematical tasks, including algebra, analytic geometry, and arithmetic. More recently, with the emergence of increasingly powerful multimodal reasoning models (Guo et al., 2025b; Gemini Team, Google, 2025), an increasing number of expert- and competition-level benchmarks (Wang et al., 2024c; He et al., 2024b; Gupta et al., 2024) have been introduced to better guide model development and evaluate their reasoning capabilities. Building on these efforts, some studies have been proposed to evaluate multimodal mathematical reasoning from model robustness to diagrammatic variations (Zhang et al., 2024e; Zou et al., 2025), reasoning with minimal or no textual context (Kamoi et al., 2024; Xiao et al., 2024), and generating structured outputs such as visual diagrams or natural language explanations (Li et al., 2025b; Park et al., 2025). The core challenge of multimodal mathematical reasoning lies in performing complex mathematical reasoning in the language space while simultaneously constructing intermediate visual content that is not explicitly provided in the input, such as drawing auxiliary lines or plotting function graphs. Therefore, success on multimodal mathematical benchmarks inherently requires dynamic reasoning on visual content, making them a crucial testbed for evaluating the “Thinking with Images” capabilities of LMMs.
@@
STEM (Science, Technology, Engineering, and Mathematics) benchmarks are designed to evaluate the reasoning capabilities of LMMs across scientific and technical domains. In this domain, early efforts focus on broad science education benchmarks, aiming to assess reasoning and knowledge across various subjects at the K–12 level (Lu et al., 2022; Hao et al., 2025; Zhou et al., 2025a; Ye et al., 2025). As the reasoning abilities of LMMs have rapidly advanced, subsequent works extend evaluations to university-level tasks that require deeper subject knowledge and complex reasoning capabilities (Yue et al., 2024a; Xia et al., 2024a; Wang et al., 2025j), and more recently, to professional-level benchmarks targeting expert reasoning in highly specialized domains (Yue et al., 2024b; Zhou et al., 2025b). Unlike the constructive but often static reasoning observed in mathematical tasks, STEM problems demand dynamic reasoning over temporal and causal phenomena, including physical motion, chemical reactions, and mechanical operations. This shift enhances the “Thinking with Images” paradigm by redefining the image as a dynamic representation of real-world processes. As a result, these benchmarks provide a testbed for evaluating a model’s understanding of physical laws and, more broadly, the sophistication of its emerging internal world model.
@@
Benchmarks for puzzles and games are designed to evaluate the capacity of LMMs for strategic planning and reasoning in dynamic, rule-governed environments. Early benchmarks introduced a range of tasks, from visual puzzles assessing logical and spatial reasoning (Zhang et al., 2024f; Ren et al., 2025; Zhang et al., 2025c; Wu et al., 2024f) to interactive games that require understanding of game mechanics and state transitions (Wang et al., 2025k; Tong et al., 2025a; Paglieri et al., 2024). More recent benchmarks have increased in complexity, incorporating dynamic video-based contexts that test physical and commonsense reasoning (Zhang et al., 2025g; Cao et al., 2024; Li et al., 2024b), as well as unstructured, competition-style puzzles demanding creative and open-ended problem solving (Wang et al., 2025l; Li et al., 2025h). Distinct from reasoning in STEM domains, the core challenge of games lies in long-horizon, prospective reasoning. This represents a form of “Thinking with Images” in which a model must visually simulate complex action sequences on an interactive game board in order to devise an optimal plan. The scalability and structured nature of these environments also makes them suitable for both training and evaluating the planning and reasoning capabilities of LMMs.
@@
Perception and reasoning benchmarks evaluate the visual capabilities of LMMs, encompassing the accurate interpretation of fine-grained visual elements and multi-step visual reasoning. One line of these tasks focuses on perceptual acuity, using high-resolution and detail-rich imagery to assess fine-grained visual recognition (Wu and Xie, 2023; Fu et al., 2024a; Wang et al., 2025m). Another line targets complex cognitive functions, including multi-step reasoning (Bi et al., 2025), object hallucination detection (Li et al., 2023a), and cognitive evaluations (Song et al., 2024; Zhao et al., 2025b; Huang et al., 2023). These benchmarks assess whether a model can move beyond superficial pattern recognition to develop a deeper, structured understanding of visual scenes that supports robust reasoning. This capability establishes the perceptual and cognitive foundation necessary for the “Thinking with Images” paradigm.
@@
Code generation from visual or textual inputs represents a key application domain of the “Thinking with Images” paradigm, requiring LMMs to generate executable programming logic based on complex visual content and corresponding instructions. Benchmarks in this area evaluate this ability across diverse contexts, such as interpreting mathematical plots and general programming-related diagrams (Li et al., 2024c; Wu et al., 2025b). One significant application is front-end development, where models are tasked with converting visual designs, ranging from hand-drawn sketches to polished UI screenshots, into functional webpage code (Si et al., 2025; Li et al., 2025i; Laurençon et al., 2024). This visual-to-code process exemplifies a distinctive form of “Thinking with Images.” It requires the model to act as a visual decompiler, parsing spatial layouts and design hierarchies to reconstruct a visual artifact into its underlying structural logic. The ability to convert visual intent into precise, executable code represents a practically valuable application of advanced reasoning LMMs.
@@
Real-world application benchmarks evaluate the ability of LMMs to handle diverse, authentic tasks that reflect practical scenarios from daily-life contexts. Work involves comprehensive, multi-task benchmarks that use authentic materials like examination questions to assess broad competency across a wide range of subjects (Zhu et al., 2024; Zhang et al., 2023c, 2025h; Chen et al., 2024b). In parallel, specialized benchmarks probe performance on specific daily-life skills, such as temporal reasoning with clocks and calendars, or visual numerical reasoning and counting (Weng et al., 2025b). The core challenge of these benchmarks is to apply reasoning to noisy, unstructured visual data by leveraging broad commonsense and world knowledge. This serves as the ultimate test for the “Thinking with Images” paradigm, assessing whether the LMMs’ visual reasoning capabilities can generalize from controlled settings to practical, real-world utility.